OpenStudy (anonymous):

A fire is sighted due west of lookout A. The bearing of the fire from lookout B, 8.2 miles due south of A, is N 33.90°W (meaning 33.90 degrees west of due north). How far is the fire from B (to the nearest tenth of a mile)?

OpenStudy (anonymous):

on sketching u will get a right angled triangle Using Cos x = adj/hyp cos33.9 = 8.2/hyp hyp=8.2 / (cos33.9)

OpenStudy (anonymous):

9.9 miles to 1 dp

OpenStudy (anonymous):

thank you